OpenStudy (anonymous):

A pet supplier has a stock of parakeets of which 10% are blue parakeets. A pet store orders 3 parakeets from this supplier. If the supplier selects the parakeets at random, what is the chance that the pet store gets exactly one blue parakeet? Choose one answer. a. 0.003 b. 0.081 c. 0.243 d. 0.027

OpenStudy (anonymous):

\[Binomial Probability = \left(\begin{matrix}n \\ x\end{matrix}\right)p^x*(1-p)^{n-x}\] where n = number of samples = 3 x = number of successes = 1 p = probability of success = .10 therefore, plugging in numbers \[\left(\begin{matrix}3 \\ 1\end{matrix}\right).1^1*(1-.1)^{3-1}\] = .243